OpenStudy (anonymous):

Find the equation in point-slope form of the lines through point P(2,0) and are A) parallel to B) perpendicular to the line x+2y=3

OpenStudy (anonymous):

point slope...

OpenStudy (aroub):

y-y1=m(x-x1) a) y-0=-1/2(x-2)

OpenStudy (aroub):

b) y-0=2(x-2)